OpenStudy (ksaimouli):

the graph of a quadratic functions (a parabola) has x-intercepts -1 and 3 and a range consisting of all numbers or equal to 4. Determine an expression for the function

OpenStudy (anonymous):

or equal means >= or <=?

OpenStudy (ksaimouli):

less than or equal to 4

OpenStudy (anonymous):

x-intercepts are the roots of function so y=(x+1)(x-3)=x^2-2x-3

OpenStudy (anonymous):

x^2-2x-3=x^2-2x+1-4=(x-1)^2-4 so f(x)=-((x-1)^2-4)=-x^2+2x+3

OpenStudy (anonymous):

the - sign added to equation for the "less than or equal to 4"

OpenStudy (anonymous):